Marine scientists categorize signature whistles of bottlenose dolphins by

typel—type ​a, type​ b, type​ c, etc. In one study of a sample of 185 whistles emitted from bottlenose dolphins in​ captivity, 97 were categorized as type a whistles.

a. Estimate the true proportion of bottlenose dolphin signature whistles that are type a whistles. Use a 99% confidence interval.The 99% confidence interval is

​(_______,_______). ​(Round to the nearest hundredth as​ needed.)

b. Interpret the result. Choose the correct answer below.

A.We are 1​% confident that the true proportion of bottlenose dolphin signature whistles that are type a whistles is in interval found in previous step.

B.We are 99% confident that the true proportion of bottlenose dolphin signature whistles that are type a whistles is ininterval found in previous step.

C.We are 99​% confident that the true proportion of bottlenose dolphin signature whistles that are type a whistles is not in interval found in previous step.

D.We are​ 100% confident that the true proportion of bottlenose dolphin signature whistles that are type a whistles is in interval found in previous step.
Answer #1

Solution: Given that n = 185 , X = 97

p = X/n = 97/185 = 0.52

q = 1-p = 1-0.52 = 0.48

99% confidence interval for the Z = 2.576

a) 99% confidence interval for the true proportion : p +/- Z*sqrt(pq/n)
= 0.52 +/- 2.576*sqrt(0.52*0.48/185)
= 0.43 , 0.61

b) => option B.